OpenStudy (azionne):

9(x - 6) = 8(x + 6) will medal

pooja195 (pooja195):



pooja195 (pooja195):

9 x X =? 9 x -6=? 8 x X = ? 8 x 6= ?

OpenStudy (pawanyadav):

You have to take 9 and 8 inside the bracket first

pooja195 (pooja195):

^also known as distributing

OpenStudy (azionne):

ok then what

pooja195 (pooja195):

\[\huge~\rm~9x-54+54=8x+48+54\] 48+54=?

OpenStudy (pawanyadav):

So 9x-54=8x+48 So9x-8x=48+54 x=102

OpenStudy (anonymous):

9(x-6)=8(xt6) 9x-54=8x=48 9x-8x=54+48 x=102

pooja195 (pooja195):

There was no need to give the answer let the asker put in the effort to get the answer rather than sppon feed the answer.

OpenStudy (anonymous):

Solve for x by simplifying both sides of the equation, then isolating the variable. x=102

OpenStudy (anonymous):

lol. But I wanted to. Just mind your own business ;) :*